This says to use systems of equations to solve: I need help. Thanks!

There is an initiation fee to join the Pine River Country Club, as well as monthly dues. The total cost after 7 months' membership will be $3025, and after 18 mos, $3850. Find both the initiation fee and the monthly dues.

Answer by stanbon(48552)

There is an initiation fee to join the Pine River Country Club, as well as monthly dues. The total cost after 7 months' membership will be $3025, and after 18 mos, $3850. Find both the initiation fee and the monthly dues.
0 solutions
---------------------
Let "d" be the dues per month and i be the initiation fee:
Cost(7 months) = i+7d=3025
Cost(18 months)= i+18d=3850
1st: i+7d=3025
2nd: i+18d=3850
Subtract 1st from 2nd to get:
3rd: 11d=825
dues= $75.00 per month
Substitue this value of d into 1st to solve for "i":
i+7(75)=3025
i= $2500.00 (this is the initiation fee)
Cheers,
Stan H.